Write an equation of the line below.
x y
(0,4)
(5,7)

Answer by stanbon(75887)

Write an equation of the line below.
x y
(0,4)
(5,7)
--------
slope = (7-4)/(4-0) = 3/4
------------------------------
Equation Form: y = mx + b
You know m = 3/4 and y = 4 when x = 0
------------------------------
Solve for m,x,y and solve for "b":
4 = (3/4)*0 = b
b = 4
------------------
Equation:
y = (3/4)x +4
=====================
Cheers,
Stan H.
=====================